Demand in a market dominated by two firms (a Cournot duopoly) is determined according to: P = 300 – 4(Q1 + Q2), where P is the market price, Q1 is the quantity demanded by Firm 1, and Q2 is the quantity demanded by Firm 2. The marginal cost and average cost for each firm is constant; AC=MC = $74. The cournot-duopoly equilibrium profit for each firm is
Answer #1

Cournot quantity for each firm is q1 = q2 = (a - c)/3b = (300 - 71)/3*4 = 19.08333 units

Cournot price = P = 300 - 4*(19.08333 + 19.08333) = 147.3333

Profit for each firm = (P - AC)*q1 = (147.3333 - 71)*19.08333

= $1456.69

Hence the answer is $1456.69